OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (hba):

b^2-4ac i guess :)

OpenStudy (anonymous):

b^2 - 4ac definitely. No need for guessing here. This is called the discriminant and will be the portion under the radical. If it is positive and gives a perfect square, then you can solve by factoring.

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

but i also have to solve it can u help me there also??

OpenStudy (hba):

Here a=2 b=7 and c=3

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]That is your solution for "x" for equations of the form: ax^2 + bx + c = 0 which is what you have: 2x^2 + 7x + 3 = 0 so, a=2, b=7, and c=3 So, just substitute, and yes, this one can be solved by factoring, but you can also use the quadratic formula as presented here.

OpenStudy (anonymous):

How's this going for you now with this detailed explanation?

OpenStudy (anonymous):

Having a hard time getting started?

OpenStudy (anonymous):

srry was afk

OpenStudy (anonymous):

-7+/7^2 -4*2*3 ------------- 2 * 2 -7 +/ 49 + 16 ------------ 4

OpenStudy (anonymous):

-7-/7^2 -4*2*3 ------------- 2 * 2 -7 -/ 49 + 16 ------------ 4

OpenStudy (anonymous):

are those right? so far

OpenStudy (anonymous):

No, arithmetic errors: -4*2*3 is not equal to 16

OpenStudy (anonymous):

Think of -4*2*3 as (-4)*(2)*(3) and (-a)(b) = -ab example: (-6) * (8) = -(6) * (8) = -48

OpenStudy (anonymous):

sorry im confused because ive done these before i just kinda forgot but when u showed me the first way to solve them i remember but this last part confused me

OpenStudy (anonymous):

Your confusion should be cleared up if you read my posts and do the substitutions and do the work. Start with my corrections to your work. If you really are having trouble multiplying a negative number by a positive number, then I cannot help you. I could give you the final answer, but that will not help you in the slightest. Not at all. The goal is for you to be able to do these on your own and you now have all the explanation I can give you. Good luck.

OpenStudy (anonymous):

btw can u read this that way when i get the answer u can help me with a way of putting my answer in.

OpenStudy (anonymous):

For Fill in the blank items, you may type in any digits between 0 and 9, – for a negative sign, / for a fraction bar, or a decimal point. Do NOT type in commas, dollar signs, percent signs, variables, or spaces. If there are extra spaces at the end of your answer, the computer will mark that problem incorrect

OpenStudy (anonymous):

so the -4 * 2 * 3 would be -24?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so -7+/7^2 -4*2*3 ------------- 2 * 2 -7 +/ 49 + 24 ------------ 4 --------------------------------------------- -7-/7^2 -4*2*3 ------------- 2 * 2 -7 -/ 49 + 16 ------------ 4

OpenStudy (anonymous):

-7-/7^2 -4*2*3 ------------- 2 * 2 -7 -/ 49 + 24 ------------ 4

OpenStudy (anonymous):

7^2 -4*2*3 is not equal to 49 + 24 You already corrected your work on -4 * 2 * 3 to be equal to -24

OpenStudy (anonymous):

srry i see now that i forgot the negative

OpenStudy (anonymous):

np. just make the correction and continue. You're starting to get this now. You're doing fine, just keep up the good work.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 49 + -24 = 25 25 SR = 5 -7 + 5 = -2 -- = 0.5 4

OpenStudy (anonymous):

-2/4 is not equal to 0.5

OpenStudy (anonymous):

and 49 + -24 = 25 25 sr = 5 -7 - 5 = -12 -- = -3 4

OpenStudy (anonymous):

srry -2

OpenStudy (anonymous):

-3 is correct, but 0.5 is not. So, you have half of your answer so far.

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

(x + 1/2)(x + 3) = (2x + 1)(x + 3) = 0 = 2x^2 + 7x + 3

OpenStudy (anonymous):

That's your factored answer.

OpenStudy (anonymous):

0.o

OpenStudy (anonymous):

You're welcome. Good luck to you in all of your studies.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thx for the recognition! @Vince34

OpenStudy (anonymous):

np

OpenStudy (anonymous):

daang 3 years ago